# Joint estimation of a two-phase spin rotation beyond classical limit

### v1

@@ II.2 Influence of crosstalk between the pump and the signal
The first three steps of the measurement scheme split the pump mode $a_{1,0}$ into two halves and establish a $\pi/2$ SU(2) rotation between $a_{1,1}$ and $a_{1,-1}$. In the ideal scenario, the signals are not affected by pumps as their polarizations are orthogonal. The unwanted couplings are, however, inevitable between sublevels of $F=1$ and $F=2$, hence they introduce noise into the signal modes. Such imperfection degrades the simultaneously measured squeezing in both subspaces, and whose influence can be simulated by adding Gaussian distributed noise to the signal modes according to
@@
where $\eta_{S}$ and $\eta_{A}$ represent complex sources of white noise satisfying $\left\langle\eta_{S}\right\rangle=\left\langle\eta_{A}\right\rangle=0$ and $\left\langle\eta_{S}^{2}\right\rangle=\left\langle\eta_{A}^{2}\right\rangle=1$. The noise strength for $a_{S}$ and $a_{A}$ are assumed to be equal $\delta\eta_{S}=\delta\eta_{A}=\delta\eta$ for simplicity here.
To calibrate the strength of the noise $\delta\eta$, we perform numerical calculations to assess the fluctuations of the simultaneously measured $Q_{yz}$ and $Q_{xz}$ for polar state, which remains unaffected by pump phase noise. Figure S1 illustrates a comparison between the experimental results and our numerical simulations, which reveals a value of approximately $\delta\eta=0.3$. It is worth pointing out that our numerical simulations incorporate a calibrated detection noise of ${\Delta(N_{2,\pm 1}^{\text{out}}-N_{1,\pm 1}^{\text{out}}})=24$ in the atomic population difference.
Figure S1: Calibration of the signal-pump crosstalk noise strength with polar state. The blue solid line represents the numerical simulations of $\Delta Q_{xz}^{\mathrm{meas}}$ as a function of the crosstalk noise strength, while the orange color dashed line corresponds to that of $\Delta Q_{yz}^{\mathrm{meas}}$. The experimental results are depicted by the blue and orange dots, which align reasonably close with the corresponding numerical simulations at $\delta\eta\approx 0.3$.
II.3 Influence of pump phase noise
When measuring squeezing in both subspaces, the phases of the two coherent pumps are required to differ by $\pi/2$. Due to the imperfection of MW pulses, fluctuation of the pump phase needs to be considered as well. While the polar state is insensitive to this noise, spin-nematic squeezed state is extremely sensitive to it, which can be modeled by adding Gaussian distributed random phase into the pump mode
@@ III.1 Sensitivity
III.2 Standard quantum limit of joint parameter estimation
@@
where the factor of 2 arises from the presence of three interference modes. In the present work, two phases are estimated simultaneously, leading to the SQL as
@@ IV.1 Initial state preparation
We prepare an initial state with about 26500 ${}^{87}$b atoms condensed in the ground $|F=1,m_{F}=0\rangle$ hyperfine state, confined in a crossed dipole trap with trapping frequencies around $2\pi\times(190,90,120)$ Hz along three orthogonal directions, by first removing atoms occupying the other states with a gradient magnetic field. The bias magnetic field is set at 0.838 G, which offers a quadratic magnetic Zeeman shift of $q_{B}\approx 2\pi\times 50$ Hz $=13|c_{2}|$ to inhibit spin-mixing. To further purify the initial polar state, two resonant microwave pulses coupling $|F=1,m_{F}=\pm 1\rangle$ to $|F=2,m_{F}=\pm 2\rangle$ are applied, followed by a probe light pulse beam to flush away the remaining |±$|\pm 1\rangle$ atoms. Subsequent spin-mixing evolution is initiated by switching on a microwave dressing field 12.18 MHz blue-detuned from the $|1,0\rangle$ to $|2,0\rangle$ transition, quenching $q$ from $13|c_{2}|$ to near $|c_{2}|$.
IV.2 Microwave pulse synthesizing
During the experiment, a sequence of phase-coherent microwave pulses with controllable $\sigma_{+/-}$-polarization is needed. It is synthesized by combining a microwave signal generator (SMB100A) with output frequency set at 6.6 GHz, and a four-channel AWG (DN2.663-04), output set at 200 MHz, using a Mini-Circuits mixer. The power stabilization is achieved by a power feedback loop. The polarization control is achieved by adjusting the relative phase between the two microwave antennas, which are placed along different spatial directions, hence carrying different polarization components. The signals of the two antennas are generated from two AWG channels, thus they remain phase coherent.
@@ IV.3 Calibration of atom number detection
The detection signal comes from absorption imaging, therefore the detected atom number accuracy is of great importance to the measurement. We prepare initial polar states with all atoms in $|F=1,m_{F}=0\rangle$ at different numbers, then apply a resonant RF $\pi/2$ pulse to generate coherent spin state $(|+1\rangle+|-1\rangle)^{\otimes N}/2^{N/2}$ and measure the variance of number difference between $|1,\pm 1\rangle$ states. As shown in  Fig. S3, we find the number difference variance matches with that of the coherent state projection noise, which affirms the detection accuracy.
Figure S3: Calibration of the atom number detection accuracy. The solid line denotes $\sqrt{N}$, the theoretical number difference variance of the coherent state. The blue dots show the measured variance, which match well with the solid line, affirming the accuracy of atom number detection.
@@ IV.4 RF amplitude calibration
The small angle RF rotation is carried out by two resonant RF pulses with $\pi$ phase difference, and the rotation angle is adjusted by changing pulse duration. We prepare the coherent state $(-i|+1\rangle+\sqrt{2}|0\rangle-i|-1\rangle)^{\otimes N}/2^{N}$ with $\langle\rho_{0}\rangle=1/2$ by RF rotation applied to the initial polar state. Subsequently, a small-angle RF rotation is applied along the same axis, resulting in a change in $\langle\rho_{0}\rangle$ proportional to the rotation angle $\alpha$, expressed as $\langle\rho_{0}\rangle=1/2+\alpha$ for small values of $\alpha$. The population changes with different pulse duration, providing a linear calibration to the rotation angle, as shown in Fig. S4.
Figure S4: The calibration of RF rotation angle. We use two resonant RF pulses with $\pi$ phase difference to realize small angle rotation. The rotation angle is adjusted by fixing one pulse and changing the duration of the other. The blue dots show the measured $\langle\rho_{0}\rangle$ changes for coherent state $(-i|+1\rangle+\sqrt{2}|0\rangle-i|-1\rangle)^{\otimes N}/2^{N}$ under such rotations, and the solid line depicts a linear fit of the measured data.

### v2

@@ II.3 Influence of crosstalk between “pump” and “signal”
The first three steps of the measurement scheme split the pump mode $a_{1,0}$ into two halves and establish a $\pi/2$ SU(2) rotation between $a_{1,1}$ and $a_{1,-1}$. In the ideal situation, the signals are not affected by pumps as their polarizations are orthogonal. Unwanted couplings are, however, inevitable between sublevels of $F=1$ and $F=2$, and they introduce noise into the signal modes. Such imperfections impair the simultaneously measured squeezing in both subspaces, whose influence can be simulated by adding Gaussian distributed noise to the signal modes according to
@@
where $\eta_{S}$ and $\eta_{A}$ represent complex sources of white noise satisfying $\left\langle\eta_{S}\right\rangle=\left\langle\eta_{A}\right\rangle=0$. The noise strength for $a_{S}$ and $a_{A}$ are assumed to be equal $\delta\eta_{S}=\delta\eta_{A}=\delta\eta$ for simplicity here.
To calibrate the strength of the noise $\delta\eta$, we perform numerical calculations to assess the fluctuations of the simultaneously measured $Q_{yz}$ and $Q_{xz}$ for polar state, which remains unaffected by pump phase noise. Figure S3 illustrates a comparison between the experimental results and our numerical simulations, which pins down a value of approximately $\delta\eta=0.3$. It is worth pointing out that our numerical simulations also incorporate a calibrated detection noise of ${\Delta(N_{2,\pm 1}^{\text{out}}-N_{1,\pm 1}^{\text{out}}})=24$ in the atomic population difference.
II.4 Influence of pump mode phase noise
When measuring squeezing in both spin-nematic subspaces, the phases of the two coherent pumps are required to differ by $\pi/2$. Due to the imperfection of MW pulses, fluctuation of the pump phase needs to be considered as well. While the polar state is insensitive to this noise, spin-nematic squeezed state is extremely sensitive to it, which is modeled by adding Gaussian distributed random phase into the pump mode
@@ II.4 Influence of pump mode phase noise
Figure S3: Calibration of the signal-pump crosstalk noise strength with polar state. The blue solid line represents the numerical simulations of $\Delta Q_{xz}^{\mathrm{meas}}$ as a function of the crosstalk noise strength, while the orange color dashed line corresponds to that of $\Delta Q_{yz}^{\mathrm{meas}}$. The experimental results are depicted by the blue and orange dots, which align reasonably close to the corresponding numerical simulations at $\delta\eta\approx 0.3$.
@@ III.1 Phase sensitivities
The phases to be estimated $\varphi_{1}$ and $\varphi_{2}$ are encoded in the alternatively parameterized form $\varphi_{1}=\alpha\cos\Theta$ and $\varphi_{2}=\alpha\sin\Theta$, where $\alpha$ of the imprinted phase is controlled experimentally by the duration of the RF pulse while $\Theta$ by the RF phase.
III.2 Standard quantum limit of two-phase joint estimation
@@
where the factor of 2 arises from the presence of three interference modes (or three paths) [44, 50]. In the present work, two phases are estimated simultaneously, leading to the SQL as
@@
Figure S5: Two distinct scenarios in two-parameter estiamtion. (a) $N$ atoms are divided into two halves with each group estimating a single phase. (b)Every one of the $N$ atoms are encoded with two phases simultaneously.
It should be noted that the SQL Eq. (S17) corresponds to the scenario depicted in the Fig. S5(b) below, where each atom is encoded with two phases, and is measured simultaneously, i.e., within one shot of absorption imaging. In this scenario, every atom contributes to the estimation of both parameters. In the main tex, the metrological gain is referenced to Eq. (S17), as our work corresponds to the scenario depicted in Fig. S5(b).
In the scenario shown in Fig. S5(a), $N$ atoms are divided into two halves equally, with each estimating a single phase. Consequently, the SQL for each group is given by $\Delta\varphi_{\mathrm{SQL}}=1/{2\sqrt{N/2}}$. Referenced to SQL of this scenario, the metrological gain reported in the main text will be 3 dB higher.
@@ IV.1 Initial state preparation
We prepare an initial state with about 26500 87Rb atoms condensed in the ground $|F=1,m_{F}=0\rangle$ hyperfine state, confined in a crossed dipole trap with trapping frequencies around $2\pi\times(190,90,120)$ Hz along three orthogonal directions, by first removing atoms occupying other states with a gradient magnetic field. The bias magnetic field is set at 0.838 G, which offers a quadratic magnetic Zeeman shift of $q_{B}\approx 2\pi\times 50$ Hz $=13|c_{2}|$, which is sufficiently large to inhibit spin-mixing. To further purify the initial polar state, two resonant microwave pulses coupling $|F=1,m_{F}=\pm 1\rangle$ to $|F=2,m_{F}=\pm 2\rangle$ are applied, followed by a probe light beam pulse to flush away the remaining |±1$\pm$ 1$\rangle$| ± 1 ⟩ atoms. Subsequent spin-mixing evolution is initiated by switching on a microwave dressing field 12.18 MHz blue-detuned from the |1,0|1,0$\rangle$| 1 , 0 ⟩ to |2,0|2,0$\rangle$| 2 , 0 ⟩ transition, which quenches $q$ from $13|c_{2}|$ to near $|c_{2}|$.
IV.2 MW pulse synthesizing
During the experiment, a sequence of phase-coherent microwave pulses with controllable $\sigma_{+/-}$-polarization is needed. It is synthesized by combining a microwave signal generator (SMB100A) with output frequency set at 6.6 GHz, and a four-channel AWG (DN2.663-04), output set at 200 MHz, using a Mini-Circuits mixer. The power stabilization is achieved by a power feedback loop. The polarization control is achieved by adjusting the relative phase between the two microwave antennas, which are placed along different spatial directions, hence carrying different polarization components. The signals of the two antennas are generated from two AWG channels, and they remain phase coherent.
@@ IV.3 Calibration of atom number detection
The detection signal comes from absorption imaging, therefore the detected atom number accuracy is of great importance to the measurement. We prepare initial polar states with all atoms in $|F=1,m_{F}=0\rangle$ at different numbers, then apply a resonant RF $\pi/2$ pulse to generate coherent spin state $(|+1\rangle+|-1\rangle)^{\otimes N}/2^{N/2}$ and measure the variance of number difference between |1,±1|1$\pm$ 1$\rangle$| 1 , ± 1 ⟩ states. As shown in Fig. S6, we find the number difference variance matches with that of the coherent state projection noise, which affirms the detection accuracy.
Figure S6: Calibration of the atom number detection accuracy. The solid line denotes $\sqrt{N}$, the theoretical number difference variance of a coherent state. The blue dots show the measured variance, which match well with the solid line, affirming the accuracy of atom number detection.
@@ IV.4 RF amplitude calibration
The small angle RF rotation is carried out by two resonant RF pulses with $\pi$ phase difference, and the rotation angle is adjusted by changing pulse duration. We prepare the coherent state $(-i|+1\rangle+\sqrt{2}|0\rangle-i|-1\rangle)^{\otimes N}/2^{N}$ with $\langle N_{0}\rangle/N=1/2$ by RF rotation applied to the initial polar state. Subsequently, a small-angle RF rotation is applied along the same axis, resulting in a change in $\langle N_{0}\rangle/N$ proportional to the rotation angle $\alpha$, expressed as $\langle N_{0}\rangle/N=1/2+\alpha$ for small values of $\alpha$. The population changes with different pulse duration, providing a linear calibration to the rotation angle, as shown in Fig. S7.
Figure S7: The calibration of RF rotation angle. We use two resonant RF pulses with $\pi$ phase difference to realize small angle rotation. The rotation angle is adjusted by fixing one pulse and changing the duration of the other. The blue dots show the measured $\langle N_{0}\rangle/N$ changes for coherent state $(-i|+1\rangle+\sqrt{2}|0\rangle-i|-1\rangle)^{\otimes N}/2^{N}$ under such rotations, and the solid line depicts a linear fit of the measured data.